Scoring a hole-in-one is the greatest shot a golfer can make. Once 8 professional golfers each made holes-in-one on the 7th hole at the same golf course at the same tournament. It has been found that the estimated probability of making a hole-in-one is 1/2722 for male professionals. Suppose that a sample of 8 professional male golfers is randomly selected.
What is the probability that at least one of these golfers makes a hole-in-one on the 13th hole at the same tournament?

Answer by nyc_function(2741)

"at least one" is the complementary event for "none"

So, we calculate the probability that none of them score a hole-in-one. Then we

subtract this probability from 1, because if it isn't true that no-one does,

then at least one of them does.

For each golfer, prob of not scoring hole-in-one

= 1 - 1/2722

= 2721/2722

Hence prob that all 8 fail to do so

= (2721/2722)^8

= 0.997 correct to 3 decimal places.

Therefore, the probability that at least one does

= 1 - 0.997

= 0.003